# Meta-rater: A Multi-dimensional Data Selection Method for Pre-training Language Models

### v1

@@ Acknowledgement
We express sincere thanks to InternTrain Team of Shanghai Artificial Intelligence Laboratory, especially Yang Gao, for their kind help in the experiment.
@@ Appendix A Ratings
The full list of 25 raters utilized in this study, including 11 rule-based natural language quality signals, 3 data importance scores, and 11 model-based quality scores is shown in Table 6.
@@
_terminal_punctution_mark Indicates whether a line ends with a terminal punctuation mark. A terminal punctation mark is defined as one of: ".", "!", "?", "”".
@@ B.1 Meta-rater Weights
As shown in Table 11, the learned weights of quality raters reveal interesting patterns about quality dimension importance.
  1. 1.
    Educational Value contributes most significantly (5.64%), aligning with findings from QuRating and FineWeb-Edu.
  2. 2.
    Writing Style contributes minimally (0.05%), consistent with QuRating’s finding that high Writing Style data did not surpass random baseline.
  3. 3.
    Natural language signals remain important despite simplicity, particularly those detecting non-alphabetical content.
  4. 4.
    Within PRRC, Reasoning (4.44%) and Professionalism (4.05%) contribute substantially, while Cleanliness (1.17%) shows less impact.
@@ C.1 Annotation Model
We selected 7 powerful LLMs as candidates for annotation: Qwen-2.5-72B-Instruct [^4] https://huggingface.co/Qwen/Qwen2.5-72B-Instruct, Qwen-2-72B-Instruct [^5] https://huggingface.co/Qwen/Qwen2-72B-Instruct, Llama-3.3-70B-Instruct [^6] https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct, Llama-3.1-70B-Instruct [^7] https://huggingface.co/meta-llama/Llama-3.1-70B-Instruct, Llama-3-70B-Instruct [^8] https://huggingface.co/meta-llama/Meta-Llama-3-70B-Instruct, gpt-4o [^9] https://openai.com/index/hello-gpt-4o/, and gpt-3.5-turbo-0125101010https://openai.com/index/chatgpt/. To determine the most suitable model, we constructed a validation dataset comprising 1,000 instances, drawing from a wide range of sources including Wikipedia, Books, Reddit, StackExchange, ArXiv, and CommonCrawl. These candidate LLMs, along with gpt-4111111https://openai.com/index/gpt-4/, were then used to score the validation dataset based on the prompts outlined in Appendix C.2. To assess the consistency of the candidate models relative to gpt-4, we computed the Kendall tau correlation score between their respective scores. The results, evaluated across four dimensions, consistently indicated that Llama-3.3-70B-Instruct outperformed the others, leading to its selection as our final annotation model.
@@ C.2 Prompts for Annotation
The four prompts used to evaluate Professionalism, Readability, Reasoning, and Cleanliness are presented in Figures 11, 12, 13, and 14. Using these prompts, Llama-3.3-70B-Instruct was tasked with annotating 1 million randomly sampled documents from SlimPajama, utilizing a maximum context window length of 128k tokens. After applying the necessary filtering and cleaning procedures, a total of 934,278 document-score pairs were retained. These pairs were subsequently divided into training, development, and test sets in an 8:1:1 ratio, resulting in 747,422 training pairs, 93,428 development pairs, and 93,428 test pairs. Moreover, specific examples in annotated SlimPajama are provided:
  • Examples of documents rated from 0 to 5 in terms of Professionalism are presented in Figures 15, 16, 17, 18, 19, and 20. These six documents include excerpts from a web page, a nursery rhyme, a magazine article, a popular science article, and two academic papers.
  • Examples of documents rated from 0 to 5 in terms of Readability are presented in Figures 21, 22, 23, 24, 25, and 26. These six documents consist of excerpts from one web page and five student essays.
  • Examples of documents rated from 0 to 5 in terms of Reasoning are presented in Figures 27, 28, 29, 30, 31, and 32. These six documents consist of excerpts from three web pages and three news.
  • Examples of documents rated from 0 to 5 in terms of Cleanliness are presented in Figures 33, 34, 35, 36, 37, and 38. These six documents consist of excerpts from six web pages.
@@ Appendix G Distribution of Raters
The distribution of 11 rule-based natural language quality signals is shown in Figures 4, 5, 6, and 7. The distribution of three data importance scores is shown in Figure 8. The distribution of 11 model-based quality scores is shown in Figures 9 and 10.
@@ H.4 Analysis of Proxy Models
Data Selection Method ARC-E ARC-C SciQ SIQA WG HS RACE OBQA
Random
   50B tokens 61.91 31.57 90.90 42.99 55.41 51.45 33.59 32.20
   100B tokens 66.33 33.53 92.80 43.71 59.59 57.35 34.35 36.20
Meta-rater (Ours)
   50B tokens 68.31 35.92 91.00 41.97 54.93 53.11 33.73 32.20
   100B tokens 72.10 37.54 92.90 43.91 60.14 58.99 35.12 37.00
Table 16: Full downstream tasks results of scaling experiment.

### v3

@@ Acknowledgement
This work is supported by Shanghai Artificial Intelligence Laboratory. We express sincere thanks to InternTrain Team of Shanghai Artificial Intelligence Laboratory, especially Yang Gao, for their kind help for the pre-training experiments.
@@ Appendix A Ratings
The full list of 25 raters utilized in this study, including 11 rule-based natural language quality signals, 3 data importance scores, and 11 model-based quality scores is shown in Table 5. The spearman correlation heatmap among 25 quality metrics is in Figure 4.
@@
_terminal_punctution_mark Indicates whether a line ends with a terminal punctuation mark. A terminal punctation mark is defined as one of: ".", "!", "?", """.
@@
Figure 4: The spearman correlation heatmap among 25 quality metrics.
Figure 5: The loss landscape of proxy model losses over the first two principal components derived from the PCA of 25 quality scores.
@@ C.1 Annotation Model
We selected 7 powerful LLMs as candidates for annotation: Qwen-2.5-72B-Instruct [^3] https://huggingface.co/Qwen/Qwen2.5-72B-Instruct, Qwen-2-72B-Instruct [^4] https://huggingface.co/Qwen/Qwen2-72B-Instruct, Llama-3.3-70B-Instruct [^5] https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct, Llama-3.1-70B-Instruct [^6] https://huggingface.co/meta-llama/Llama-3.1-70B-Instruct, Llama-3-70B-Instruct [^7] https://huggingface.co/meta-llama/Meta-Llama-3-70B-Instruct, gpt-4o [^8] https://openai.com/index/hello-gpt-4o/, and gpt-3.5-turbo-0125999https://openai.com/index/chatgpt/. To determine the most suitable model, we constructed a validation dataset comprising 1,000 instances, drawing from a wide range of sources including Wikipedia, Books, Reddit, StackExchange, ArXiv, and CommonCrawl. These candidate LLMs, along with gpt-4101010https://openai.com/index/gpt-4/, were then used to score the validation dataset based on the prompts outlined in Appendix C.2. To assess the consistency of the candidate models relative to gpt-4, we computed the Kendall tau correlation score between their respective scores. The results, evaluated across four dimensions, consistently indicated that Llama-3.3-70B-Instruct outperformed the others, leading to its selection as our final annotation model.
@@ C.2 Prompts for Annotation
The four prompts used to evaluate Professionalism, Readability, Reasoning, and Cleanliness are presented in Figures 13, 14, 15, and 16. Using these prompts, Llama-3.3-70B-Instruct was tasked with annotating 1 million randomly sampled documents from SlimPajama, utilizing a maximum context window length of 128k tokens. After applying the necessary filtering and cleaning procedures, a total of 934,278 document-score pairs were retained. These pairs were subsequently divided into training, development, and test sets in an 8:1:1 ratio, resulting in 747,422 training pairs, 93,428 development pairs, and 93,428 test pairs. Moreover, specific examples in annotated SlimPajama are provided:
  • Examples of documents rated from 0 to 5 in terms of Professionalism are presented in Figures 17, 18, 19, 20, 21, and 22. These six documents include excerpts from a web page, a nursery rhyme, a magazine article, a popular science article, and two academic papers.
  • Examples of documents rated from 0 to 5 in terms of Readability are presented in Figures 23, 24, 25, 26, 27, and 28. These six documents consist of excerpts from one web page and five student essays.
  • Examples of documents rated from 0 to 5 in terms of Reasoning are presented in Figures 29, 30, 31, 32, 33, and 34. These six documents consist of excerpts from three web pages and three news.
  • Examples of documents rated from 0 to 5 in terms of Cleanliness are presented in Figures 35, 36, 37, 38, 39, and 40. These six documents consist of excerpts from six web pages.
@@ Appendix G Distribution of Raters
The distribution of 11 rule-based natural language quality signals is shown in Figures 6, 7, 8, and 9. The distribution of three data importance scores is shown in Figure 10. The distribution of 11 model-based quality scores is shown in Figures 11 and 12.
@@ H.4 Analysis of Proxy Models
Model Size Method ARC-E ARC-C SciQ SIQA WG HS RACE OBQA
3.3B Random 66.33 33.53 92.80 43.71 59.59 57.35 34.35 36.20
Meta-rater 72.10 37.54 92.90 43.91 60.14 58.99 35.12 37.00
7.2B Random 67.77 36.43 91.10 42.73 60.29 53.02 34.73 37.00
Meta-rater 71.34 39.76 92.80 44.32 60.45 58.97 36.08 38.20
Table 15: Full downstream tasks results of 3.3B and 7.2B models.
@@ H.5 Analysis of Data Domain
Appendix I Evalution Results on MMLU and NaturalQuestions
We also evaluate pre-trained models on two challenging tasks, with results shown in Table 19. Our analysis reveals several important insights:
1. Scale limitations: For MMLU, all 1.3B models perform near random-chance level (25%), confirming previous findings that smaller models struggle with this benchmark. This aligns with observations in prior work Liu et al. (2025); Wettig et al. (2024) showing that models below 7B parameters typically perform at or slightly above random chance on MMLU regardless of training methodology.
2. Consistent patterns: Despite the overall low performance, Meta-rater still shows a slight improvement over random selection in NaturalQuestions for both model scales (2.30% vs. 2.13% for 1.3B; 6.87% vs. 6.28% for 3.3B). This suggests our method’s benefits extend to knowledge-intensive tasks, though the absolute performance remains limited by model capacity.
3. Scaling effects: The significant jump in NaturalQuestions performance from 1.3B to 7.2B models (approximately 5x improvement) indicates that model scale is particularly important for knowledge-intensive tasks. This is consistent with the literature showing that knowledge retrieval capabilities improve non-linearly with model size.
Model Method MMLU NQ
1.3B Random 25.99 2.13
QuRating-Educational Value 26.73 2.05
Meta-rater 25.89 2.30
3.3B Random 25.48 6.28
Meta-rater 26.21 6.87
7.2B Random 26.21 10.89
Meta-rater 26.24 10.42
Table 19: Downstream tasks results on challenging tasks.